Find the probability that the income-tax official will catch 3 income-tax turns with illegitimate deductions. if he randomly selects 5 returns from among 12 returns of which 6 contain illegitimate deductions


1
Expert's answer
2021-07-19T05:49:59-0400

Let "X=" the number of income-tax turns with illegitimate deductions: "X\sim Bin(n, p)."

Given "n=5, p=6\/12=0.5, q=1-p=1-0.5=0.5."


"P(X=3)=\dbinom{5}{3}(0.5)^3(0.5)^{5-3}=0.3125"

The probability that the income-tax official will catch 3 income-tax turns with illegitimate deductions is 0.3125.